### v1

# A Complete Set of Quadratic Constraints For Repeated ReLU

Abstract
This paper derives a complete set of quadratic constraints (QCs) for the repeated ReLU. The complete set of QCs is described by a collection of $2^{n_{v}}$ matrix copositivity conditions where $n_{v}$ is the dimension of the repeated ReLU. We also show that only two functions satisfy all QCs in our complete set: the repeated ReLU and a repeated ”flipped” ReLU. Thus our complete set of QCs bounds the repeated ReLU as tight as possible up to the sign invariance inherent in quadratic forms. We derive a similar complete set of incremental QCs for repeated ReLU, which can potentially lead to less conservative Lipschitz bounds for ReLU networks than the standard LipSDP approach. Finally, we illustrate the use of the complete set of QCs to assess stability and performance for recurrent neural networks with ReLU activation functions. The stability/performance condition combines Lyapunov/dissipativity theory with the QCs for repeated ReLU. A numerical implementation is given and demonstrated via a simple example.
I Introduction
This paper considers the use of quadratic constraints (QCs) for a nonlinear function. QCs are inequalities expressed as quadratic forms on the input/output graph of the function. These constraints are useful as they can be easily incorporated into stability and performance conditions for dynamical systems. This is a special case of the more general integral quadratic constraint (IQC) framework [1, 2, 3, 4]. Moreover, the related class of incremental QCs can be used to compute Lipschitz bounds on static functions.
This paper focuses on QCs and incremental QCs for the specific case of a repeated Rectified Linear Unit (ReLU). This special case is motivated the use of the scalar ReLU as the activation function in neural networks (NNs) and recurrent neural networks (RNNs). The scalar ReLU $\phi:\mathbb{R}\to\mathbb{R}$ is defined formally in Section II-B. It satisfies a number of useful properties including $\phi(v)\geq 0$ (positivity), $\phi(v)\geq v$ (positive complement), $\phi(v)(v-\phi(v))=0$ (complementarity), and slope restrictions. These properties have been used to derive QCs for the repeated ReLU [5, 6, 7, 8, 9].
QCs can be combined with Lyapunov/dissipativity theory [10, 11, 12, 13] to derive stability and performance conditions for RNNs. Related work along these lines for discrete-time RNNs is given in [14, 15, 16, 5, 8, 9, 17]. In addition, incremental QCs have been used to compute Lipschitz bounds on NNs [18]. There is a rapidly growing literature on the application of incremental QCs including extensions for $\ell_{\infty}$ perturbations [19], deep equilibrium NNs [20], and the use for designing Lipschitz networks [21, 22, 23, 24]. There is also recent work on scalability to ImageNet [25] as well as extensions to convolutional structures [26, 27] and activation functions such as MaxMin and GroupSort [28].
Our contributions to this existing literature are as follows. First, we provide the complete class of QCs satisfied by the repeated ReLU ($\mathcal{M}_{c}$ in Theorem 1). This complete set of QCs is described by a collection of $2^{n_{v}}$ matrix copositivity conditions where $n_{v}$ is the dimension of the repeated ReLU. We also show that only two functions satisfy all QCs in our complete set: the repeated ReLU and a repeated ”flipped” ReLU (Theorem 2). Thus our complete set of QCs bounds the repeated ReLU as tight as possible up to the sign invariance inherent in quadratic forms. Next we focus on incremental QCs. We derive the complete set of incremental QCs for repeated ReLU ($\mathcal{M}^{inc}$ in Theorem 3). We also derive a new subclass of incremental QCs for repeated ReLU ($\mathcal{M}^{inc}_{2}$ in Lemma 4). This subclass may be useful for Lipschitz analysis problems with large $n_{v}$ and hence the complete class cannot be used due to its computational cost. Finally, Section IV provides an application of the QCs to analyze dynamic systems with ReLU activation functions. This section also discusses aspects of the numerical implementation. A simple example illustrating the approach is given in Section V.
II Background
II-A Notation
This section introduces basic notation regarding vectors, matrices, and signals. Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the sets of real $n\times 1$ vectors and $n\times m$ matrices, respectively. Moreover, $\mathbb{R}_{\geq 0}^{n}$ and $\mathbb{R}_{\geq 0}^{n\times m}$ denote vectors and matrices of the given dimensions with non-negative entries. Thus $\mathbb{R}_{\geq 0}^{n}$ corresponds to the non-negative orthant. We let $e_{i}\in\mathbb{R}^{n_{v}}$ denote the $i^{th}$ basis vector, i.e. every entry of $e_{i}$ is zero except entry $i$ is equal to 1.
Next, define the following sets of matrices:
  • $\mathbb{S}^{n}$ is the set of real symmetric, $n\times n$ matrices.
  • $D^{n}$ is the set of $n\times n$, real diagonal matrices.
  • $COP^{n}$ is the set of $n\times n$, real, symmetric co-positive matrices. Specifically, if $M=M^{\top}\in COP^{n}$ then $x^{\top}Mx\geq 0$ for all $x\in\mathbb{R}^{n}_{\geq 0}$.
  • $DH^{n}$ is the set of $n\times n$, real, doubly hyperdominant matrices. Such matrices have non-positive off-diagonal elements while both the row sums and column sums are non-negative.
  • $\mathbb{N}$ is the set of non-negative integers.
For $M\in\mathbb{S}^{n}$, we use $M\succeq 0$ and $M\succ 0$ to denote that it is positive semidefinite or positive definite, respectively.
Finally, let $v:\mathbb{N}\to\mathbb{R}^{n}$ and $w:\mathbb{N}\to\mathbb{R}^{n}$ be real, vector-valued sequences. Define the inner product $\langle v,w\rangle:=\sum_{k=0}^{\infty}v(k)^{\top}w(k)$. The set $\ell_{2}^{n}$ is an inner product space with sequences $v$ that satisfy $\langle v,v\rangle<\infty$. The corresponding norm is $\|v\|_{2}:=\sqrt{\langle v,v\rangle}$.
II-B Quadratic Constraints for Repeated ReLU
The Rectified Linear Unit (ReLU) is a function $\phi:\mathbb{R}\to\mathbb{R}_{\geq 0}$ defined by:
$\phi(v)=\left\{\begin{array}[]{ll}0&\mbox{if }v<0\\ v&\mbox{if }v\geq 0\end{array}\right..$ (3)
The graph of the ReLU is shown in Figure 1. The repeated ReLU is the function $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ defined by $\Phi(v)=(\phi(v_{1}),\,\phi(v_{2}),\ldots,\,\phi(v_{n_{v}}))$
$v$$w=\phi(v)$
Figure 1: Graph of scalar ReLU $\phi$.
This paper mainly focuses on quadratic constraints (QCs) satisfied by the repeated ReLU. The notion of a QC is formally defined next.
Definition 1
A function $F:\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{v}}$ satisfies the Quadratic Constraint (QC) defined by $M\in\mathbb{S}^{2n_{v}}$ if the following inequality holds for all $v\in\mathbb{R}^{n_{v}}$ and $w=F(v)$:
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M\begin{bmatrix}v\\ w\end{bmatrix}\geq 0.$ (4)
QCs are, in general, conservative bounds on the graph of a function in the sense that a given vector pair $(v,\,w)$ may satisfy (4) and yet $w=F(v)$ may not hold. However, these bounds are useful as they can be easily incorporated into stability and performance conditions for dynamical systems. One application of QCs to stability analysis will be given in Section IV.
QCs for repeated ReLU can be derived using simple properties of the scalar ReLU. Specifically, the scalar ReLU satisfies the following useful properties that have been used previously in the literature [5, 6, 7, 8, 9]:
  1. 1.
    Positivity: The scalar ReLU is non-negative for all inputs: $\phi(v)\geq 0$ $\forall v\in\mathbb{R}$.
  2. 2.
    Positive Complement: The scalar ReLU satisfies $\phi(v)\geq v$ $\forall v\in\mathbb{R}$.
  3. 3.
    Complementarity: The graph of the scalar ReLU is identically on the line of slope 0 (when $v\leq 0$) or the line of slope 1 (when $v\geq 0$). Thus, $\phi(v)\,(v-\phi(v))=0$ $\forall v\in\mathbb{R}$.
  4. 4.
    Positive Homogeneity: The scalar ReLU is homogeneous for all non-negative constants: $\phi(\beta v)=\beta\phi(v)$ $\forall v\in\mathbb{R}$ and $\forall\beta\in\mathbb{R}_{\geq 0}$.
The next lemma gives two QCs for the repeated ReLU based on these properties. Both QCs are known in the literature and the second one, defined by $\mathcal{M}_{2}$ below, is a restatement of Theorem 2 in [9].
Lemma 1
The repeated ReLU $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ satisfies the QCs defined by any matrices in the following sets:
$\mathcal{M}_{1}$ :=$:=\left\{\begin{bmatrix}0&Q_{1}\\ Q_{1}&-2Q_{1}\end{bmatrix}\,:\,Q_{1}\in D^{n_{v}}\right\},$ (5)
$\mathcal{M}_{2}$ :=$:=\left\{\begin{bmatrix}-I&I\\ 0&I\end{bmatrix}^{\top}Q_{2}\begin{bmatrix}-I&I\\ 0&I\end{bmatrix}\,:\,Q_{2}\in COP^{2n_{v}}\right\}.$ (6)
Proof:
Take any $M_{1}\in\mathcal{M}_{1}$. Complementary of scalar ReLU implies that for all $v\in\mathbb{R}^{n_{v}}$ and $w=\Phi(v)$,
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M_{1}\begin{bmatrix}v\\ w\end{bmatrix}=2\sum_{i=1}^{n_{v}}(Q_{1})_{ii}w_{i}(v_{i}-w_{i})=0.$
Thus $\Phi$ satisfies the QC defined by any $M_{1}\in\mathcal{M}_{1}$.
Next, take any $M_{2}\in\mathcal{M}_{2}$. The positivity and positive complement properties of scalar ReLU imply that for all $v\in\mathbb{R}^{n_{v}}$ and $w=\Phi(v)$,
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M_{2}\begin{bmatrix}v\\ w\end{bmatrix}=\begin{bmatrix}w-v\\ w\end{bmatrix}^{\top}Q_{2}\begin{bmatrix}w-v\\ w\end{bmatrix}\geq 0.$
The inequality holds because $Q_{2}$ is copositive and $\left[\begin{smallmatrix}w-v\\ w\end{smallmatrix}\right]$ is elementwise nonnegative. Thus $\Phi$ satisfies the QC defined by any $M_{2}\in\mathcal{M}_{2}$. ∎
Repeated ReLU also satisfies QCs defined for the more general class of slope restricted nonlinearities. Specifically, scalar ReLU is slope restricted to $[0,1]$, i.e.:
$0\leq\frac{\phi(v_{2})-\phi(v_{1})}{v_{2}-v_{1}}\leq 1,\,\,\forall v_{1},\,v_{2}\in\mathbb{R},\,v_{1}\neq v_{2}.$
The next lemma defines a set of QCs for repeated ReLU based on this slope-restriction. These are a special case of Zames-Falb multipliers [29]. The lemma follows from existing results for slope restricted nonlinearities (Section 3.5 of [30] and [31] [^1] The results in these references are stated for repeated monotone nonlinearities. A similar fact holds for nonlinearities with slope restricted to $[0,1]$ by a transformation of the input-output data.).
Lemma 2
The repeated ReLU $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ satisfies the QCs defined by any matrix in the following set:
$\mathcal{M}_{3}:=\left\{\begin{bmatrix}0&Q_{3}^{\top}\\ Q_{3}&-(Q_{3}+Q_{3}^{\top})\end{bmatrix}\,:\,Q_{3}\in DH^{n_{v}}\right\}.$ (7)
The next section gives an independent proof that repeated ReLU satisfies all QCs defined by $\mathcal{M}_{3}$.
III Main Results
III-A Complete Set of QCs for Repeated ReLU
The previous section introduced three classes of QCs for repeated ReLU. There are other QCs for repeated ReLU in the literature, e.g. polytopic QCs [9, 8]. This raises the following question: What is the largest class of QCs satisfied by the repeated ReLU? This question is addressed by our main results in this subsection.
Let $D_{\pm 1}^{n}\subset D^{n}$ denote the set of $n\times n$ diagonal matrices where each diagonal entry is ±$\pm 1$. There are $2^{n_{v}}$ matrices in this set. Next, let matrices $M\in\mathbb{S}^{2n_{v}}$ and $D\in D_{\pm 1}^{n_{v}}$ be given. Let $M_{D}\in\mathbb{S}^{n_{v}}$ denote the following matrix:
$M_{D}:=\begin{bmatrix}D\\ \frac{1}{2}(I+D)\end{bmatrix}^{\top}M\begin{bmatrix}D\\ \frac{1}{2}(I+D)\end{bmatrix}.$ (8)
Given this notation, define the following set of matrices:
$\mathcal{M}_{c}:=\left\{M\in\mathbb{S}^{2n_{v}}\,:\,M_{D}\in COP^{n_{v}}\,\,\forall D\in D_{\pm 1}^{n_{v}}\right\}.$ (9)
The matrix dimension $2n_{v}$ is omitted from the notation $\mathcal{M}_{c}$ but should be clear from context. Each matrix in $\mathcal{M}_{c}$ satisfies a collection of $2^{n_{v}}$ copositivity conditions; one condition for each $D\in D_{\pm 1}^{n_{v}}$. The next theorem states that $\mathcal{M}_{c}$ defines the complete class of QCs for the repeated ReLU.
Theorem 1
The repeated ReLU $\Phi$ satisfies a QC defined by $M\in\mathbb{S}^{2n_{v}}$ if and only if $M\in\mathcal{M}_{c}$.
Proof:
($\Leftarrow$) Suppose $M\in\mathcal{M}_{c}$. Take any $v\in\mathbb{R}^{n_{v}}$ and let $w=\Phi(v)$. Define a diagonal matrix $\bar{D}\in D_{\pm 1}^{n_{v}}$ and vector $\bar{v}\in\mathbb{R}^{n_{v}}_{\geq 0}$ by [^2] Here we use the convention that sign():{1,+$(\cdot):\mathbb{R}\to\{-1,+1\}$ with sign$(x)=+1$ if $x\geq 0$ and sign$(x)=-1$ if $x<0$.:
$\bar{D}_{kk}:=\mbox{sign}(v_{k})\mbox{ and }\bar{v}_{k}:=|v_{k}|.$ (10)
Thus $v=\bar{D}\bar{v}$ and $w=\frac{1}{2}(I+\bar{D})\bar{v}$. The equality for $w$ follows because $\frac{1}{2}(1+\bar{D}_{kk})\bar{v}_{k}$ is equal to 0 when $v_{k}<0$ and is equal to $v_{k}$ when $v_{k}\geq 0$.
Let $M_{\bar{D}}$ denote the scaled matrix defined as in (8) but with $\bar{D}\in D_{\pm 1}^{n_{v}}$. Thus $M_{\bar{D}}$ is copositive since $M\in\mathcal{M}_{c}$ by assumption. Combining these facts gives:
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M\begin{bmatrix}v\\ w\end{bmatrix}=\bar{v}^{\top}M_{\bar{D}}\bar{v}\geq 0.$ (11)
Since $v$ was arbitrary the above holds for all $v\in\mathbb{R}^{n_{v}}$; namely, the repeated ReLU $\Phi$ satisfies the QC defined by $M\in\mathcal{M}_{c}$ as required.
($\Rightarrow$) This direction is by contrapositive. Assume $M\notin\mathcal{M}_{c}$. Then $M_{D}$ in (8) is not copositive for some $D\in D_{\pm 1}^{n_{v}}$. Thus there exists $\bar{v}\in\mathbb{R}_{\geq 0}^{n_{v}}$ such that $\bar{v}^{\top}M_{D}\bar{v}<0$. Define $v:=D\bar{v}$ and $w:=\frac{1}{2}(I+D)\bar{v}$. If $D_{kk}=-1$ then $v_{k}\leq 0$ and $w_{k}=0$. If $D_{kk}=+1$ then $v_{k}\geq 0$ and $w_{k}=v_{k}$. Hence $w=\Phi(v)$ and
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M\begin{bmatrix}v\\ w\end{bmatrix}=\bar{v}^{\top}M_{D}\bar{v}<0.$ (12)
Thus repeated ReLU does not satisfy the QC defined by $M$.
$\mathcal{M}_{c}$ defines the complete set of QCs satisfied by repeated ReLU by Theorem 1. Section II-B discussed several existing QCs for repeated ReLU. The next lemma and corollary provide an independent proof for these existing QCs relying on the characterization in Theorem 1.
Lemma 3
The sets $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{3}$ are subsets of $\mathcal{M}_{c}$.
Proof:
The proof is given in the appendix. ∎
Corollary 1
The repeated ReLU $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ satisfies the QCs defined by any matrices in $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\mathcal{M}_{3}$.
Proof:
This follows from Theorem 1 and Lemma 3. ∎
III-B Functions Characterized by Complete Set of QCs
Theorem 1 leads to a related question: Are there other functions, besides repeated ReLU, that satisfy all QCs defined by matrices in $\mathcal{M}_{c}$? The answer to this question is yes. To show this, define a new function $\Psi:\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{v}}$ by $\Psi(v):=-\Phi(-v)$. The pair $(v,w)$ lies on the graph of $\Phi$ if and only if $(-v,-w)$ lies on the graph of $\Psi$. Therefore we call $\Psi$ the repeated flipped-ReLU. Next note that quadratic forms are even functions:
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M\begin{bmatrix}v\\ w\end{bmatrix}=\begin{bmatrix}-v\\ -w\end{bmatrix}^{\top}M\begin{bmatrix}-v\\ -w\end{bmatrix}.$ (13)
Thus $\Phi$ satisfies the QC defined by some $M$ if and only if $\Psi$ also satisfies the QC defined by $M$. The next theorem states that repeated ReLU and flipped-ReLU are the only functions that satisfy all QCs defined by $\mathcal{M}_{c}$. In other words, the set of QCs defined by $\mathcal{M}_{c}$ is as tight as possible for repeated ReLU up to the sign-invariance inherent in quadratic forms.
Theorem 2
A function $F:\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{v}}$ satisfies all QCs defined by $\mathcal{M}_{c}$ if and only if $F$ is either $\Phi$ or $\Psi$.
Proof:
($\Leftarrow$) Assume $F$ is either $\Phi$ or $\Psi$. If $F=\Phi$ then Theorem 1 implies that $F$ satisfies all QCs defined by $\mathcal{M}_{c}$. The same is true when $F=\Psi$ because quadratic forms are even functions.
($\Rightarrow$) Assume $F$ satisfies all QCs defined by $\mathcal{M}_{c}$. By Lemma 3, $\mathcal{M}_{c}$ contains any $M_{1}:=\left[\begin{smallmatrix}0&Q_{1}\\ Q_{1}&-2Q_{1}\end{smallmatrix}\right]$ with $Q_{1}$ diagonal. Select $Q_{1}$ to have all zero entries except for entry $(i,i)$. Then all $v\in\mathbb{R}^{n_{v}}$, $w=F(v)$ satisfy
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M_{1}\begin{bmatrix}v\\ w\end{bmatrix}=2(Q_{1})_{ii}w_{i}(v_{i}-w_{i})\geq 0.$ (14)
$(Q_{1})_{ii}$ can be chosen to be positive or negative. Thus (14) implies that $w_{i}=0$ or $w_{i}=v_{i}$. In other words, $F_{i}(v)$ only depends on $v_{i}$ and $F_{i}(v_{i})\in\{0,v_{i}\}$ for all $v_{i}\in\mathbb{R}$.
Lemma 3 also implies that $\mathcal{M}_{c}$ contains any $M_{2}\in\mathcal{M}_{2}$. Define $M_{2}\in\mathcal{M}_{2}$ by
$M_{2}:=\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]^{\top}Q_{2}\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]\mbox{ with }Q_{2}=\left[\begin{smallmatrix}0&0\\ 0&e_{i}e_{j}^{\top}\end{smallmatrix}\right],\,i\neq j.$ (15)
Then all $v\in\mathbb{R}^{n_{v}}$, $w=F(v)$ satisfy
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M_{2}\begin{bmatrix}v\\ w\end{bmatrix}=2w_{i}w_{j}\geq 0.$ (16)
This condition is violated if $F_{i}(v_{i})$ and $F_{j}(v_{j})$ have opposite sign for some $v_{i},v_{j}\in\mathbb{R}$. This implies: (i) $F_{i}$ is globally nonnegative for all $i=1,\ldots n_{v}$, or (ii) $F_{i}$ is globally nonpositive for all $i=1,\ldots n_{v}$.
First consider case (i). Define $M_{2}\in\mathcal{M}_{2}$ with $Q_{2}=\left[\begin{smallmatrix}e_{i}e_{j}^{\top}&0\\ 0&0\end{smallmatrix}\right]$ for some $i\neq j$. Then all $v\in\mathbb{R}^{n_{v}}$, $w=F(v)$ satisfy
$\begin{bmatrix}v\\ w\end{bmatrix}^{\top}M_{2}\begin{bmatrix}v\\ w\end{bmatrix}=2(w_{i}-v_{i})(w_{j}-v_{j})\geq 0.$ (17)
Equation (17) is violated if $F_{i}(v_{i})-v_{i}$ and $F_{j}(v_{j})-v_{j}$ have opposite sign for some $v_{i},v_{j}\in\mathbb{R}$. This implies: (iii) $F_{i}(v_{i})\geq v_{i}$ globally for all $i=1,\ldots n_{v}$, or (iv) $F_{i}(v_{i})\leq v_{i}$ globally for all $i=1,\ldots n_{v}$. Note that (iv) contradicts (i). Thus if (i) holds then each $F_{i}$ must satisfy: $F_{i}(v_{i})\in\{0,v_{i}\}$, (i) $F_{i}$ is globally nonnegative, and (iii) $F_{i}(v_{i})\geq v_{i}$ globally. Therefore each $F_{i}$ is the scalar ReLU and $F$ is the repeated ReLU, i.e. $F=\Phi$.
Next consider case (ii) above. Note that case (iii) contradicts (ii). Thus if (ii) holds then each $F_{i}$ must satisfy: $F_{i}(v_{i})\in\{0,v_{i}\}$, (ii) $F_{i}$ is globally nonpositive, and (iv) $F_{i}(v_{i})\leq v_{i}$ globally. Therefore each $F_{i}$ is the scalar flipped ReLU and $F$ is the repeated flipped ReLU, i.e. $F=\Psi$. ∎
III-C Incremental QCs
Incremental QCs, formally defined below, are a related class of quadratic constraints. Incremental QCs can be used to compute Lipschitz bounds on Neural Networks (NNs), e.g. as in LipSDP [18] and related work [19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
Definition 2
A function $F:\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{v}}$ satisfies the incremental QC defined by $M\in\mathbb{S}^{2n_{v}}$ if the following inequality holds for all $\bar{v},\,\hat{v}\in\mathbb{R}^{n_{v}}$ and $\bar{w}=F(\bar{v})$, $\hat{w}=F(\hat{v})$:
$\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}^{\top}M\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}\geq 0.$ (18)
Incremental QCs for repeated ReLU can be derived from properties of scalar ReLU. First, scalar ReLU is slope restricted in $[0,1]$. This slope restriction implies that for any $\bar{v}$, $\hat{v}\in\mathbb{R}$ and $\bar{w}=\phi(\bar{v})$, $\hat{w}=\phi(\hat{v})$:
$(\bar{w}-\hat{w})\cdot\left[(\bar{v}-\hat{v})-(\bar{w}-\hat{w})\right]\geq 0.$ (19)
It is well known that the above property can be used to show that the repeated ReLU $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ satisfies the incremental QCs defined by any matrices in the following set [18, 26, 27] :
$\mathcal{M}^{inc}_{1}$ :=$:=\left\{\begin{bmatrix}0&T_{1}\\ T_{1}&-2T_{1}\end{bmatrix}\,:\,T_{1}\in D^{n_{v}},\,T_{1}\succeq 0\right\}.$ (20)
To see this, take any $M_{1}\in\mathcal{M}^{inc}_{1}$. Then for any $\bar{v},\,\hat{v}\in\mathbb{R}^{n_{v}}$ and $\bar{w}=\Phi(\bar{v})$, $\hat{w}=\Phi(\hat{v})$,
$\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}^{\top}M_{1}\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}$
$=2\sum_{i=1}^{n_{v}}(T_{1})_{ii}\cdot(\bar{w}_{i}-\hat{w}_{i})\cdot\left[(\bar{v}_{i}-\hat{v}_{i})-(\bar{w}_{i}-\hat{w}_{i})\right].$
Each term of the sum is nonnegative due to $T_{1}\succeq 0$ and the $[0,1]$ slope restriction of scalar ReLU in (19). Thus $\Phi$ satisfies all incremental QCs defined by $\mathcal{M}^{inc}_{1}$. The above incremental QC is standard, and has been widely used in LipSDP and related work [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
Interestingly, scalar ReLU satisfies another useful property. If $\bar{v}-\hat{v}\geq 0$ then $\bar{w}-\hat{w}\in[0,\bar{v}-\hat{v}]$ while if $\bar{v}-\hat{v}\leq 0$ then $\bar{w}-\hat{w}\in[\bar{v}-\hat{v},0]$. It follows that if $(\bar{v}_{i},\hat{v}_{i},\bar{w}_{i},\hat{w}_{i})$ for $i=1,2$ are obtained from scalar ReLU then: [^3] To simplify notation let $dv_{i}:=\bar{v}_{i}-\hat{v}_{i}$ and $dw_{i}:=\bar{w}_{i}-\hat{w}_{i}$. If $dv_{1}dv_{2}\geq 0$ then $dw_{1}dw_{2}\geq 0$ and thus $(dv_{1}-dv_{2})^{2}+2dw_{1}dw_{2}\geq 0$. If $dv_{1}dv_{2}\leq 0$ then $dw_{1}dw_{2}\geq dv_{1}dv_{2}$ This implies $(dv_{1}-dv_{2})^{2}+2dw_{1}dw_{2}\geq dv_{1}^{2}+dv_{2}^{2}\geq 0$.
$\left[(\bar{v}_{1}-\hat{v}_{1})-(\bar{v}_{2}-\hat{v}_{2})\right]^{2}+2(\bar{w}_{1}-\hat{w}_{1})\cdot(\bar{w}_{2}-\hat{w}_{2})\geq 0.$ (21)
The next lemma gives a new incremental QC for the repeated ReLU based on this property.
Lemma 4
The repeated ReLU $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ satisfies the incremental QCs defined by any matrix in the following set:
$\mathcal{M}^{inc}_{2}$ :=$:=\left\{\begin{bmatrix}T_{2}&0\\ 0&S_{2}-T_{2}\end{bmatrix}\,:\,T_{2}=T_{2}^{\top}\in DH^{n_{v}},\,\right.$ (22)
$\left.S_{2}\in D^{n_{v}},\ (S_{2})_{ii}=(T_{2})_{ii},\,\sum_{k=1}^{n_{v}}T_{ik}=0\mbox{ for }i=1,\ldots,n_{v}\right\}.$
Proof:
Take any $M_{2}\in\mathcal{M}^{inc}_{2}$. The upper left block $T_{2}$ is a symmetric, doubly hyperdominant matrix. Moreover, the rows (and columns) sum to zero, i.e. $T_{2}$ has zero excess. Every symmetric, doubly hyperdominant matrix with zero excess can be decomposed as:
$T_{2}=\sum_{i,j=1}^{n_{v}}\lambda_{i,j}(e_{i}-e_{j})(e_{i}-e_{j})^{\top}.$ (23)
[…]
$M_{2}=\sum_{i,j=1}^{n_{v}}\lambda_{ij}\left[\begin{smallmatrix}(e_{i}-e_{j})(e_{i}-e_{j})^{\top}&0\\ 0&e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}\end{smallmatrix}\right].$ (24)
Then for any $\bar{v},\,\hat{v}\in\mathbb{R}^{n_{v}}$ and $\bar{w}=\Phi(\bar{v})$, $\hat{w}=\Phi(\hat{v})$,
$\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}^{\top}M_{2}\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}$
$=\sum_{i,j=1}^{n_{v}}\lambda_{ij}\left[\left((\bar{v}_{i}-\hat{v}_{i})-(\bar{v}_{j}-\hat{v}_{j})\right)^{2}+2(\bar{w}_{i}-\hat{w}_{i})\cdot(\bar{w}_{j}-\hat{w}_{j})\right]$
Each term of this sum is nonnegative by property (21) of the scalar ReLU and $\lambda_{ij}\geq 0$. Thus $\Phi$ satisfies all incremental QCs defined by $\mathcal{M}^{inc}_{2}$. ∎
III-D Complete Set of Incremental QCs for Repeated ReLU
We can build on our previous results to define the largest class of incremental QCs for repeated ReLU. Let matrices $M\in\mathbb{S}^{2n_{v}}$ and $D_{1}$, $D_{2}\in D_{\pm 1}^{n_{v}}$ be given. Let $M_{D_{1},D_{2}}\in\mathbb{S}^{2n_{v}}$ denote the following matrix:
$M_{D_{1}D_{2}}:=\left[\begin{smallmatrix}D_{1}&-D_{2}\\ \frac{1}{2}(I+D_{1})&-\frac{1}{2}(I+D_{2})\end{smallmatrix}\right]^{\top}M\left[\begin{smallmatrix}D_{1}&-D_{2}\\ \frac{1}{2}(I+D_{1})&-\frac{1}{2}(I+D_{2})\end{smallmatrix}\right].$ (25)
Given this notation, define the following set of matrices:
$\mathcal{M}^{inc}$ :=$:=\left\{M\in\mathbb{S}^{2n_{v}}\,:\,M_{D_{1}D_{2}}\in COP^{2n_{v}}\,\,\forall D_{1},\,D_{2}\in D_{\pm 1}^{n_{v}}\right\}.$ (26)
Note that $\mathcal{M}^{inc}$ involves a collection $2^{2n_{v}}=4^{n_{v}}$ conditions; one condition for each pair $D_{1}$, $D_{2}\in D^{n_{v}}_{\pm 1}$. The next theorem states that $\mathcal{M}^{inc}$ defines the complete class of incremental QCs for repeated ReLU.
Theorem 3
The repeated ReLU $\Phi$ satisfies an incremental QC defined by $M\in\mathbb{S}^{2n_{v}}$ if and only if $M\in\mathcal{M}^{inc}$.
Proof:
The proof follows by showing an incremental QC on $\Phi$ is equivalent to a structured (non-incremental) QC on a repeated ReLU with twice the dimension. Specifically, the quadratic form in the definition of an incremental QC can be equivalently written as:
$\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}^{\top}M\begin{bmatrix}\bar{v}-\hat{v}\\ \bar{w}-\hat{w}\end{bmatrix}=\begin{bmatrix}\bar{v}\\ \hat{v}\\ \bar{w}\\ \hat{w}\end{bmatrix}^{\top}R^{\top}MR\begin{bmatrix}\bar{v}\\ \hat{v}\\ \bar{w}\\ \hat{w}\end{bmatrix}$
[…]
Thus $M$ defines an incremental constraint for a repeated ReLU of dimension $n_{v}$ if and only if $R^{\top}MR$ defines a normal (non-incremental) QC for a repeated ReLU of dimension $2n_{v}$. By Theorem 1, this is equivalent to the following condition:
$\begin{bmatrix}D\\ \frac{1}{2}(I+D)\end{bmatrix}^{\top}(R^{\top}MR)\begin{bmatrix}D\\ \frac{1}{2}(I+D)\end{bmatrix}\in COP^{2n_{v}}\,\,\forall D\in D^{2n_{v}}_{\pm 1}.$ (27)
Block partition $D=\left[\begin{smallmatrix}D_{1}&0\\ 0&D_{2}\end{smallmatrix}\right]$. Then (27) simplifies, by direct multiplication, to $M_{D_{1}D_{2}}\in COP^{2n_{v}}$ for all $D_{1}$, $D_{2}\in D^{n_{v}}_{\pm 1}$. In summary, $M$ defines an incremental QC for repeated ReLU if and only if $M\in\mathcal{M}^{inc}$. ∎
Repeated ReLU satisfies the incremental QCs defined by the sets $\mathcal{M}^{inc}_{1}$ and $\mathcal{M}^{inc}_{2}$. It follows from Theorem 3 that $\mathcal{M}^{inc}_{1}$ and $\mathcal{M}^{inc}_{2}$ are subsets of $\mathcal{M}^{inc}$. In fact, verifying $\mathcal{M}^{inc}_{1}\subset\mathcal{M}^{inc}$ can be done via simple algebraic manipulations similar to those given for the proof of Lemma 3 in the Appendix. Therefore, Theorem 3 provides a unified tool for obtaining incremental QCs for the repeated ReLU.
IV Stability Analysis With ReLU QCs
IV-A Dynamic Systems With ReLU Activation Functions
Consider the interconnection shown in Figure 2 with a repeated ReLU $\Phi$ wrapped in feedback around the top channels of a nominal system $G$. This interconnection is denoted as $F_{U}(G,\Phi)$. The nominal part $G$ is a discrete-time, linear time-invariant (LTI) system described by the following state-space model:
$\begin{split}&x(k+1)=A\,x(k)+B_{1}\,w(k)+B_{2}\,d(k)\\ &v(k)=C_{1}\,x(k)+D_{11}\,w(k)+D_{12}\,d(k)\\ &e(k)=C_{2}\,x(k)+D_{21}\,w(k)+D_{22}\,d(k),\end{split}$ (28)
where $x\in\mathbb{R}^{n_{x}}$ is the state. The inputs are $w\in\mathbb{R}^{n_{w}}$ and $d\in\mathbb{R}^{n_{d}}$ while $v\in\mathbb{R}^{n_{v}}$ and $e\in\mathbb{R}^{n_{e}}$ are outputs. The interconnection $F_{U}(G,\Phi)$ is known as a linear fractional transformation (LFT) in the robust control literature [34]. The interconnection has its roots in the Lurye decomposition used in the absolute stability problem [13]. Recurrent Neural Networks (RNNs) with ReLU activation functions can be decomposed into this form.
$G$$d$$e$$\begin{smallmatrix}\phi&&\\ &\,\,\ddots&\\ &&\phi\end{smallmatrix}$$\Phi$$v$$w$
Figure 2: Interconnection $F_{U}(G,\Phi)$ of a nominal discrete-time LTI system $G$ and repeated ReLU $\Phi$.
This feedback interconnection involves an implicit equation if $D_{11}\neq 0$. Specifically, the second equation in (28) combined with $w(k)=\Phi(v(k))$ yields:
$v(k)=C_{1}\,x(k)+D_{11}\,\Phi(v(k))+D_{12}\,d(k).$ (29)
This equation is well-posed if there exists a unique solution $v(k)$ for all values of $x(k)$ and $d(k)$. Well-posedness of this equation implies that the dynamic system $F_{U}(G,\Phi)$ is well-posed in the following sense: for all initial conditions $x(0)\in\mathbb{R}^{n_{x}}$ and inputs $d\in\ell_{2}^{n_{d}}$ there exists unique solutions $x$, $v$, $w$ and $e$ in $\ell_{2}$ to the system $F_{U}(G,\Phi)$. There are simple sufficient conditions for well-posedness of (29), e.g. Lemma 1 in [5] (which relies on results in [35, 36]). Thus, we’ll assume well-posed for simplicity in the remainder of the paper.
A well-posed interconnection $F_{U}(G,\Phi)$ is internally stable if $x(k)\to 0$ from any initial condition $x(0)$ with $d(k)=0$ for $k\in\mathbb{N}$. In other words, $F_{U}(G,\Phi)$ is internally stable if $x=0$ is a globally asymptotically stable equilibrium point with no external input. A well-posed interconnection $F_{U}(G,\Phi)$ has finite induced-$\ell_{2}$ gain if there exists $\gamma<\infty$ such that the output $e$ generated by any $d\in\ell_{2}^{n_{d}}$ with $x(0)=0$ satisfies $\|e\|_{2}\leq\gamma\,\|d\|_{2}$. This bound is denoted as $\|F_{U}(G,\Phi)\|_{2\to 2}\leq\gamma$.
IV-B Stability and Performance Condition
We next state a stability and performance condition for $F_{U}(G,\Phi)$ using the ReLU QCs derived in the previous section.
Theorem 4
Consider the system $F_{U}(G,\Phi)$ with the LTI system $G$ defined in (28) and $\Phi:\mathbb{R}^{n_{v}}\to\mathbb{R}_{\geq 0}^{n_{v}}$ a repeated ReLU. Assume $F_{U}(G,\Phi)$ is well-posed. Also assume there exists a $2n_{v}\times 2n_{v}$ matrix $M\in\mathcal{M}_{c}$, scalar $\gamma>0$, and $P\in\mathbb{S}^{n_{x}}$ with $P\succeq 0$ such that $LMI(P,M,\gamma^{2})\prec 0$ where:
$LMI(P,M,\gamma^{2}):=\begin{bmatrix}A^{\top}PA-P&A^{\top}PB_{1}&A^{\top}PB_{2}\\ B_{1}^{\top}PA&B_{1}^{\top}PB_{1}&B_{1}^{\top}PB_{2}\\ B_{2}^{\top}PA&B_{2}^{\top}PB_{1}&B_{2}^{\top}PB_{2}-\gamma^{2}I\end{bmatrix}$
$\hskip 21.68121pt+\begin{bmatrix}C_{2}^{\top}\\ D_{21}^{\top}\\ D_{22}^{\top}\end{bmatrix}\begin{bmatrix}C_{2}^{\top}\\ D_{21}^{\top}\\ D_{22}^{\top}\end{bmatrix}^{\top}+\begin{bmatrix}C_{1}^{\top}&0\\ D_{11}^{\top}&I\\ D_{12}^{\top}&0\end{bmatrix}M\begin{bmatrix}C_{1}^{\top}&0\\ D_{11}^{\top}&I\\ D_{12}^{\top}&0\end{bmatrix}^{\top}.$ (30)
Then $F_{U}(G,\Phi)$ is internally stable and $\|F_{U}(G,\Phi)\|_{2\to 2}<\gamma$.
Proof:
This theorem is a standard dissipation result [12, 10, 11, 13, 37]. A proof, similar to the one given in [17], is given for completeness.
The LMI is strictly feasible by assumption and hence it remains feasible under small perturbations: $LMI(P+\epsilon I,M,\gamma^{2})+\epsilon I\prec 0$ for some sufficiently small $\epsilon>0$. The system $F_{U}(G,\Phi)$ is well-posed by assumption. Hence it has a unique causal solution $(x,w,v,e)$ (all in $\ell_{2}$) for any given initial condition $x(0)$ and input $d\in\ell_{2}^{n_{d}}$.
Define a storage function by $V\left(x\right):=x^{\top}(P+\epsilon I)x$. Left and right multiply the perturbed LMI by $[x(k)^{\top},\,w(k)^{\top},\,d(k)^{\top}]$ and its transpose. The result, applying the dynamics (28), gives the following condition:
$V\left(x(k+1)\right)-V\left(x(k)\right)-\gamma^{2}d(k)^{\top}d(k)+e(k)^{\top}e(k)$
$+\begin{bmatrix}v(k)\\ w(k)\end{bmatrix}^{\top}M\begin{bmatrix}v(k)\\ w(k)\end{bmatrix}+\epsilon\left[\begin{smallmatrix}x(k)\\ w(k)\\ d(k)\end{smallmatrix}\right]^{\top}\left[\begin{smallmatrix}x(k)\\ w(k)\\ d(k)\end{smallmatrix}\right]\leq 0.$
$M\in\mathcal{M}_{c}$ defines a valid QC for repeated ReLU by Theorem 1. Hence the term involving $M$ is non-negative. Thus the dissipation inequality simplifies to:
$\begin{split}&V\left(x(k+1)\right)-V\left(x(k)\right)+e(k)^{\top}e(k)\\ &\leq(\gamma^{2}-\epsilon)d(k)^{\top}d(k)-\epsilon x(k)^{\top}x(k).\end{split}$ (31)
Internal stability and the $\ell_{2}$ gain bound for $F_{U}(G,\Phi)$ follow from this inequality. Specifically, if $d(k)=0$ for all $k$ then (31) implies the following Lyapunov inequality:
$V(x(k+1))-V(x(k))\leq-\epsilon x(k)^{\top}x(k).$
Hence $V$ is a Lyapunov function and the system is globally asymptotically stable (Theorem 27 in Section 5.9 of [38]).
Next, assume $x(0)=0$ and $d\in\ell_{2}^{n_{d}}$. Summing (31) from $k=0$ to $k=T-1$ and using $V(x(0))=0$ yields:
$V\left(x(T)\right)+\sum_{k=0}^{T-1}e(k)^{\top}e(k)\leq\sum_{k=0}^{T-1}(\gamma^{2}-\epsilon)d(k)^{\top}d(k).$
Note that $V(x(T))\geq 0$ because $P$ is positive semidefinite. Moreover, the right side is upper bounded by $(\gamma^{2}-\epsilon)\|d\|_{2}^{2}$ for all $T\in\mathbb{N}$. This implies that $e\in\ell_{2}$ and $\|e\|_{2}<\gamma\|d\|_{2}$. ∎
IV-C Numerical Implementation Using $\mathcal{M}_{c}$
Let $\{D_{1},\ldots,D_{2^{n_{v}}}\}$ denote the $2^{n_{v}}$ entries of $D_{\pm 1}^{n_{v}}$. Define the following convex optimization based on Theorem 4:
$\begin{split}&\min_{P=P^{\top},\,M=M^{\top},\,\gamma^{2}}\gamma^{2}\\ &P\succeq 0,\,\,\,LMI(P,M,\gamma^{2})\prec 0,\\ &\begin{bmatrix}D_{i}\\ \frac{1}{2}(I+D_{i})\end{bmatrix}^{\top}M\begin{bmatrix}D_{i}\\ \frac{1}{2}(I+D_{i})\end{bmatrix}\in COP^{n_{v}},\,\,i=1,\ldots 2^{n_{v}}.\end{split}$ (32)
If the optimization is feasible then $F_{U}(G,\Phi)$ is stable. Moreover, the optimization returns the tightest (smallest) upper bound on the $\ell_{2}$ gain (using a quadratic storage and the complete set of ReLU QCs). The set of copositive matrices is a closed, convex cone (Proposition 1.24 in [39]). It follows that $\mathcal{M}_{c}$ is a convex cone and (32) is a convex optimization.
However, simply testing if a matrix is not copositive is an NP-complete problem [40]. Thus it is common to use relaxations for copositivity conditions. One simple sufficient condition is: if $S=S^{\top}$ is positive semidefinite and $N=N^{\top}$ is elementwise nonnegative then $S+N$ is copositive (Remark 1.10 in [39]). This relaxation is exact for $n_{v}\leq 4$ [41] but not for $n_{v}\geq 5$ (Example 1.30 in [39]). We can use this relaxation to reformulate (32) as a semidefinite program (SDP):
$\begin{split}&\min_{P=P^{\top},\,M=M^{\top},\,\gamma^{2},N_{1},\ldots,N_{2^{n_{v}}}}\gamma^{2}\\ &P\succeq 0,\,\,\,LMI(P,M,\gamma^{2})\prec 0,\\ &\begin{bmatrix}D_{i}\\ \frac{1}{2}(I+D_{i})\end{bmatrix}^{\top}M\begin{bmatrix}D_{i}\\ \frac{1}{2}(I+D_{i})\end{bmatrix}-N_{i}\succ 0,\,\,i=1,\ldots 2^{n_{v}}\\ &N_{i}\in\mathbb{R}^{n_{v}}_{\geq 0},\,\,i=1,\ldots 2^{n_{v}}\end{split}$ (33)
This SDP has $2^{n_{v}}$ positive semidefiniteness constraints arising from our relaxation of the copositivity constraints in $\mathcal{M}_{c}$. This is in addition to the elementwise nonnegativity constraints on $N_{i}$ for $i=1,\ldots,2^{n_{v}}$. This limits the use of $\mathcal{M}_{c}$ to problems where $n_{v}$ is relatively small. Larger values of $n_{v}$ will require the use of QCs that are subsets of $\mathcal{M}_{c}$.
IV-D Numerical Implementation With Subsets of $\mathcal{M}_{c}$
One useful subset of $\mathcal{M}_{c}$ consists of all combinations the existing QCs described in Section II-B. These correspond to QCs defined by matrices in the following set:
$\mathcal{M}_{123}$ :=$:=\left\{M_{1}+M_{2}+M_{3}\,:\,M_{i}\in\mathcal{M}_{i},\,i=1,2,3\right\}.$ (34)
The set $\mathcal{M}_{c}$ is a convex cone (as noted above) and $\mathcal{M}_{i}\subset\mathcal{M}_{c}$ ($i=1,2,3$) by Lemma 3. It follows that $\mathcal{M}_{123}\subset\mathcal{M}_{c}$.
Recall that Theorem 2 states that only repeated ReLU and flipped ReLU satisfy all constraints in the complete set $\mathcal{M}_{c}$. It is interesting that the proof only uses the QCs defined by matrices in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The proof does not require the use of Zames-Falb QCs defined by $\mathcal{M}_{3}$. Define another subset $\mathcal{M}_{c}$ but without the Zames-Falb QCs:
$\mathcal{M}_{12}$ :=$:=\left\{M_{1}+M_{2}\,:\,M_{i}\in\mathcal{M}_{i},\,i=1,2\right\}.$ (35)
The next result states that the Zames-Falb QCs defined by $\mathcal{M}_{3}$ do not increase the class of QCs when combined with $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
Theorem 5
The sets $\mathcal{M}_{12}$ in (35) and $\mathcal{M}_{123}$ in (34) are equal, i.e. $\mathcal{M}_{12}=\mathcal{M}_{123}$.
Proof:
($\subseteq$) Consider any $M=M_{1}+M_{2}\in\mathcal{M}_{12}$. Then $M=M_{1}+M_{2}+M_{3}\in\mathcal{M}_{123}$. with $M_{3}:=0\in\mathcal{M}_{3}$.
($\supseteq$) Consider any $M=M_{1}+M_{2}+M_{3}\in\mathcal{M}_{12}$. There exists $Q_{1}\in D^{n_{v}}$, $Q_{2}\in COP^{2n_{v}}$ and $Q_{3}\in DH^{n_{v}}$ such that
$M=\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]^{\top}Q_{2}\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]+\left[\begin{smallmatrix}0&(Q_{1}+Q_{3})^{\top}\\ (Q_{1}+Q_{3})&-(2Q_{1}+Q_{3}+Q_{3}^{\top})\end{smallmatrix}\right].$
Decompose $Q_{3}=D-P$ where $D\in D^{n_{v}}$ with $D_{ii}=(Q_{3})_{ii}$. All elements of $P=D-Q_{3}$ are nonnegative because $Q_{3}$ is doubly hyperdominant. Next define $\hat{Q}_{1}:=Q_{1}+D\in D^{n_{v}}$ and $\hat{Q}_{2}:=Q_{2}+\left[\begin{smallmatrix}0&P^{\top}\\ P&0\end{smallmatrix}\right]$. The term $\left[\begin{smallmatrix}0&P^{\top}\\ P&0\end{smallmatrix}\right]$ is copositive as it is symmetric and elementwise nonnegative (Remark 1.10 of [39]). Hence $\hat{Q}_{2}$ is copositive. Finally, define $\hat{M}_{1}\in\mathcal{M}_{1}$ and $\hat{M}_{2}\in\mathcal{M}_{2}$ corresponding to $\hat{Q}_{1}$ and $\hat{Q}_{2}$, respectively. It can be verified directly that $M=\hat{M}_{1}+\hat{M}_{2}\in\mathcal{M}_{12}$. ∎
The stability/performance condition in Theorem 4 with $\mathcal{M}_{12}$ leads to the following SDP:
$\begin{split}&\min_{P=P^{\top},\,M=M^{\top},\,Q_{1},Q_{2},N_{2},\gamma^{2}}\gamma^{2}\\ &P\succeq 0,\,\,\,LMI(P,M,\gamma^{2})\prec 0,\\ &M=\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]^{\top}Q_{2}\left[\begin{smallmatrix}-I&I\\ 0&I\end{smallmatrix}\right]+\left[\begin{smallmatrix}0&Q_{1}\\ Q_{1}&-2Q_{1}\end{smallmatrix}\right]\\ &Q_{1}\in D^{n_{v}},\,Q_{2}-N_{2}\succeq 0,\,N_{2}\in\mathbb{R}_{\geq 0}^{2n_{v}\times 2n_{v}}.\end{split}$ (36)
Here we have used the relaxation described above for the constraint $Q_{2}\in COP^{2n_{v}}$. Adding the Zames-Falb QCs defined by $\mathcal{M}_{3}$ will not reduce the conservatism in this SDP. [^5] One minor point is that Theorem 5 states that $\mathcal{M}_{12}=\mathcal{M}_{123}$ where $\mathcal{M}_{2}$ has a copositivity condition. The same set equality holds if we define $\mathcal{M}_{2}$ using the relaxation for the copositivity condition as in SDP (36). The SDP (36) will be more computationally efficient, but also more conservative, than the SDP (33) formulated using $\mathcal{M}_{c}$. In fact, we provide a numerical example in Section V where the use of $\mathcal{M}_{c}$ provides improved results compared to the use of $\mathcal{M}_{12}$. This numerical example indicates that $\mathcal{M}_{12}$ is a strict subset of the complete set $\mathcal{M}_{c}$. Thus $\mathcal{M}_{c}$ should be used when possible as it is the largest possible class of QCs for repeated ReLU.
V Example
We consider the interconnection $F_{U}(G,\Phi)$ shown in Figure 2 with a repeated ReLU $\Phi$ wrapped in feedback around the top channels of a nominal system $G$. The nominal part $G$ is a discrete-time, LTI system (28) with the following data:
$A$ $:=\left[\begin{smallmatrix}4.1819&-4.1122&4.1810&-3.4344\\ 9.5280&-9.1573&8.4496&-6.2574\\ 8.6800&-7.3880&6.0327&-4.0370\\ 2.8000&-1.7500&1.2060&-0.7209\end{smallmatrix}\right],$
$B_{1}$ $:=\alpha\times\left[\begin{smallmatrix}9.528&8.68&5.60&2.00\\ 17.36&11.20&4.00&0\\ 11.20&4.00&0&0\\ 2.00&0&0&0\end{smallmatrix}\right],\,\,B_{2}:=\left[\begin{smallmatrix}1\\ 1\\ 1\\ 1\\ 1\end{smallmatrix}\right],$
$C_{1}$ $:=\left[\begin{smallmatrix}-0.5000&0.4875&-0.2250&-0.0250\\ -0.4250&0.6500&-0.6155&0.3604\\ 0.1100&0.1282&-0.3323&0.3064\\ 0.5645&-0.5248&0.2859&-0.0793\end{smallmatrix}\right],\,\,C_{2}:=\left[\begin{smallmatrix}1&1&1&1&1\end{smallmatrix}\right],$
$D_{11}$ $:=\alpha\times\left[\begin{smallmatrix}0&0&0&0\\ -1&0&0&0\\ -0.85&-1&0&0\\ 0.22&-0.85&-1&0\end{smallmatrix}\right],\,\,D_{12}:=0,\,D_{21}:=0,\,D_{22}:=0.$
This system is a modification of Example 2 in [42] using the lifting described in [17]. The repeated ReLU $\Phi$ has dimension $n_{v}=4$ and $\alpha$ is a gain variation included on the $(v,w)$ channels associated with $\Phi$. If $\alpha=0$ then dynamics from $d$ to $e$ are governed by the ”nominal” LTI system $G_{0}$ described by $(A,B_{2},C_{2},D_{22})$. The induced gain in this case is equal to the $H_{\infty}$ norm: $\|G_{0}\|_{\infty}=39.8$.
We can use Theorem 4 to compute an upper bound on the induced $\ell_{2}$ gain of $F_{U}(G,\Phi)$ when $\alpha\geq 0$. We compute bounds using the sets of QCs described by $\mathcal{M}_{c}$ and $\mathcal{M}_{12}$ and the corresponding SDPs (33) and (36), respectively. Figure 3 shows the two bounds for 20 linearly spaced values of $\alpha$ from 00 to $0.6$. Both curves agree with the nominal gain $\|G_{0}\|_{\infty}=39.8$ at $\alpha=0$. The complete set $\mathcal{M}_{c}$ provides a less conservative bound than the QCs defined by $\mathcal{M}_{12}$. It took 17.5sec and 22.1 sec to compute all 20 points on the curves for $\mathcal{M}_{12}$ and $\mathcal{M}_{c}$, respectively. Thus the computational costs are similar for this example although we expect the computation with $\mathcal{M}_{c}$ to grow more rapidly with $n_{v}$. The curve with $\mathcal{M}_{123}$ was also computed but is not shown since it is indistinguishable from the curve for $\mathcal{M}_{12}$. This is expected since Theorem 5 states $\mathcal{M}_{12}=\mathcal{M}_{123}$.
Figure 3: Bound on induced $\ell_{2}$ gain for system $F_{U}(G,\Phi)$ vs. $\alpha$ using QCs defined by $\mathcal{M}_{c}$ and $\mathcal{M}_{12}$. The complete set $\mathcal{M}_{c}$ provides a less conservative (smaller) bound on the gain.
VI Conclusions
This paper derived a complete set of quadratic constraints (QCs) for the repeated ReLU. The complete set of QCs is described by a collection of $2^{n_{v}}$ matrix copositivity conditions where $n_{v}$ is the dimension of the repeated ReLU. The relationship between our complete set and existing QCs has been carefully discussed. We also derived a similar complete set of incremental QCs for repeated ReLU. We illustrate the use of the complete set of QCs to assess stability and performance for recurrent neural networks with ReLU activation functions. We will study, as future work, the conservatism (if any) introduced by using the complete set of repeated ReLU QCs for stability/performance analysis.
VII Acknowledgments
The authors acknowledge AFOSR Grant #FA9550-23-1-0732 for funding of this work. The authors also acknowledge Carsten Scherer for asking the question about the complete set of ReLU quadratic constraints.